OpenStudy (anonymous):

The table shows the median home prices in Florida. What is the equation of a trend line that models a relationship between time and home prices? Use the equation to predict the median home price in 2020. Year 1940 1950 1960 1970 1980 1990 2000 Median Price($) $20,200 $39,900 $50,400 $62,800 $89,100 $98,700 $119,000 A. y = 1646.67x + 20,200; $151,900 B. y = 49.4x + 20,200; $119,000 C. y = 49.4x + 20,200; $151,900 D. y = 1646.67x + 20,200; $119,000

OpenStudy (cwrw238):

x = 0 for 1940 as y = 0 + 20,200 for 1940 for 1950, x = 10 y = 36666 which is close to given price looks like A is the closest equation for 2020 y = 1646.67*80 + 20,200 = 151933

OpenStudy (anonymous):

thank you so much

OpenStudy (cwrw238):

yw